an author of mystery books received a royalty check for $350,000 and decides to keep $50,000 and invest the remainder. He invests some in certificates of desposit (CDs) that have an annual yield at 3% and the remainder in mutual funds which have an anticipated return of 4.5%. If the total returns from the two investments are $11,625 per year, find the amount in each investment?

Answer by mananth(16946)

CD's 3.00% per annum ------------- Amount invested =x
MF 4.50% per annum ------------ Amount invested = y =
300000
Interest----- 11625

CD's 3.00% per annum ---x
MF 4.50% per annum ---y
Total investment
x + 1 y= 300000 -------------1
Interest on both investments
3.00% x + 4.50% y= 11625
Multiply by 100
3 x + 4.5 y= 1162500.00 --------2
Multiply (1) by -3
we get
-3 x -3 y= -900000.00
Add this to (2)
0 x 1.5 y= 262500
divide by 1.5
y = 175000
CD's 3.00% $ 125000
MF 4.50% $ 175000

CHECK
125000 --------- 3.00% ------- 3750.00
175000 ------------- 4.50% ------- 7875.00
Total -------------------- 11625.00